OpenStudy (anonymous):

You have already invested $500 in a stock with an annual return of 11%. How much of an additional $800 should be invested at 12% and how much at 6% so that the total return on the entire $1,300 is 9%? (Round each answer to the nearest cent.)

OpenStudy (anonymous):

1300*9/100=500*11/100 + x*6/100 + (800-x)*12/100 solve to get x=566 and 2/3